OpenStudy (anonymous):

Simplify the expression. -5+i/2i Show all steps.

OpenStudy (anonymous):

\[\frac{-5+i}{2i}\] Multiplying numerator and denominator by i we find; \[\frac{-5+i}{2i} = \frac{-5+i}{2i} \times \frac{i}{i}\] \[= \frac{(-5+i) \times i}{2i \times i}= \frac{(-5i+i^2)}{2i^2}\] \[= \frac{(-5i+(-1))}{2(-1)}= \frac{-5i-1}{-2}= \frac{-(5i+1)}{-2}= \frac{1+5i}{2}\] \[= \frac{1}{2}+\frac{5}{2}i\] is the required simplified form. @Greenanole

OpenStudy (anonymous):

Reason: \[i^2 = -1\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

@Greenanole medal please.